geerky42 (geerky42):

Four feet of wire is to be used to form a square and a circle. How much of the wire should be used for the square and how much should be used for the circle to enclose the maximum total area?

OpenStudy (anonymous):

ok so

geerky42 (geerky42):

Don't help me if you don't want to. You don't have to...

OpenStudy (anonymous):

jus give me a sec dont rush the right answer

geerky42 (geerky42):

This is Calculus question, I noticed you asked a question "what is -5/6 + 1/3 " I don't think you can handle my problem. Thanks, though.

OpenStudy (anonymous):

Length for the circle = 2 pi x 0.4829 = 3.0342ft Length for the square = 4 - 3.034 = 0.9658 ft

OpenStudy (anonymous):

ohh yes and jst fyi always ask those kind of questions because i love to giv out medals

OpenStudy (anonymous):

im a nice person

geerky42 (geerky42):

Frankly, that's stupid, but ok. Your answer is wrong, by the way. I have answer key, I just don't know how to get this this answer...

OpenStudy (anonymous):

what are the answers then, he choices

geerky42 (geerky42):

All of wire should be used for the circle to enclose the maximum total area. I just don't know how to prove it.

geerky42 (geerky42):

It's not multiple choice, it's free response.

OpenStudy (anonymous):

Mmm, seem a wee bit tetchy Geerky42. The proof is simple a sircle encloses the max. area with the smallest perimeter (circumference).

geerky42 (geerky42):

I don't know if I have to show work for this question, though. I used calculus and my work said that I have to use some wire for square, I don't understand why I end up like that. Here's my work if you are interested.

geerky42 (geerky42):

Hopefully everyone can read it.

OpenStudy (anonymous):

Let; A = total area X = side of square R= radius of circle A= x^2 +pi r^2 2xdx + 2pi rdr = 0 dx/dr = -pi r/x L = 4x + 2pi r 4dx + 2pi dr = 0 dx/dr =- pi/2 -pi/2 =- pi r/x x =r/2 4 = 4(r/2) +2 pi r 4 = 2r(1+pi) r = 2/(1+pi) = 2/4.1416 = 0.4829 ft Length for the circle = 2 pi x 0.4829 = 3.0342ft Length for the square = 4 - 3.034 = 0.9658 ft

OpenStudy (anonymous):

http://malaysia.answers.yahoo.com/question/index?qid=20080720213959AAKdhat Thats my brothers answer ad page

OpenStudy (anonymous):

and page

OpenStudy (anonymous):

hello

geerky42 (geerky42):

Our answers are not the same, yet we are both wrong.

OpenStudy (anonymous):

????????????????? what grade are you in and what school do you go to

geerky42 (geerky42):

I'm junior at public high school.

OpenStudy (anonymous):

ohhhhhhhhhh

geerky42 (geerky42):

Can you help me, please? @zepdrix @hartnn @RadEn